Bims Corporation uses the weighted-average method in its process costing system. The Assembly Department started the month with 4,200 units in its beginning work in process inventory that were 70% complete with respect to conversion costs. An additional 66,500 units were transferred in from the prior department during the month to begin processing in the Assembly Department. There were 29,000 units in the ending work in process inventory of the Assembly Department that were 60% complete with respect to conversion costs.

What were the equivalent units for conversion costs in the Assembly Department for the month?
Answer #1

Under the weighted average method, equivalent units are calculated based on 2 things: units completed and transferred out and units in ending work in process inventory.

Units completed and transferred = Beginning work in process units + Units started - Ending work in process units

= 4,200 + 66,500 - 29,000

= 41,700

Equivalent units of conversion costs = units completed and transferred out + units in ending work in process inventory x Percentage complete

= 41,700 + 29,000 x 60%

= 41,700 + 17,400

= 59,100

Please ask if you have any query related to the question. Thank you